# Boundary transfer matrices arising from quantum symmetric pairs

### v1

@@ 2.5. Boundary gauge transformation and transfer matrices
Let $[{\operatorname{Mod}}_{\sf fd}(A)]$ be the Grothendieck group of the category of finite-dimensional $A$-modules, i.e., the abelian group generated by the isomorphism classes $[V]$ with $V\in{\operatorname{Mod}}_{\sf fd}(A)$ with the relations $[U]+[W]=[V]$ for every short exact sequence $0\to U\to V\to W\to 0$ in ${\operatorname{Mod}}_{\sf fd}(A)$. Since $A$ is a bialgebra, $[{\operatorname{Mod}}_{\sf fd}(A)]$ is a ring, with multiplication map given by $[V]\cdot[W]=[V\otimes W]$. Moreover, it is commutative, since $A$ is quasitriangular and thus $[V\otimes W]=[W\otimes V]$.
Given a cylindrical structure $(\psi,J,K)$ on $(A,B)$, we aim to construct a ring homomorphism from $[{\operatorname{Mod}}_{\sf fd}(A)]$ to $B$. Our approach relies on a distinguished family of gauge transformations.
@@
$[V]\mapsto\mathscr{T}^{(V)}\coloneqq\mathop{\sf Tr}_{2}\big{(}\overline{\mathbb{K}}_{\bullet,V}\big{)}$ (2.22)
where $V\in{\operatorname{Mod}}_{\sf fd}(A)$. In particular, $[\mathscr{T}^{(V)},\mathscr{T}^{(W)}]=0$ for all $V,W\in{\operatorname{Mod}}_{\sf fd}(A)$.
@@
$=\mathop{\sf Tr}_{2,3}\Big{(}\big{(}(\overline{J}_{V,W}^{-1})_{23}\cdot({\overline{{\mathbb{K}}}}_{\bullet,W})_{13}\big{)}^{\mathsf{t}_{3}}\cdot\big{(}\big{(}R_{V^{\overline{\psi}},W}\big{)}_{23}\cdot({\overline{{\mathbb{K}}}}_{\bullet,V})_{12}\big{)}^{\mathsf{t}_{3}}\Big{)}$ (2.25)
@@ 3.1. Balanced and sovereign Hopf algebras
Let $A$ be a Hopf algebra with invertible antipode $S\colon A\to A$. Then, one has
@@
The Drinfeld element is the invertible element $u\coloneqq m((S^{-1}\otimes{\operatorname{id}})(R))$, where $m\colon A\otimes A\to A$ denotes the multiplication map. It satisfies the identities [^6] The element $u$ is $u_{2}^{-1}$ in [Dri90].
@@
  • A balance is also a K-matrix. Namely, if $b$ is a balance, then $(\psi,J,K)=({\operatorname{id}},R_{21}^{-1},b)$ is a cylindrical structure on $(A,A)$, which is gauge-equivalent to $({\operatorname{id}},R,1)$ with respect to $g=b^{-1}$.
  • A half-balance [KT09, ST09] is an element $h\in A^{\times}$ such that $\Delta(h)=(h\otimes h)\cdot R$ and $h^{2}$ is central (and hence a balance). Then $(\psi,J,K)=(\operatorname{Ad}(h),1\otimes 1,h)$ is a cylindrical structure on $(A,A)$, which is gauge-equivalent to $({\operatorname{id}},R,1)$ with respect to $g=h^{-1}$. $\triangledown$
@@ 3.3. Boundary transfer matrix for balanced Hopf algebras
Let $B\subseteq A$ be a right coideal subalgebra, and $(\psi,J,K)$ a cylindrical structure on $(A,B)$.
@@
Let $g\in A^{\times}$ and $(\overline{\psi},\overline{J},\overline{K})$ the gauge-transformed cylindrical structure defined in (2.15). Then $g$ is a boundary gauge transformation if
@@
Let $A$ be a balanced Hopf algebra, $B\subseteq A$ a right coideal subalgebra, and $(\psi,J,K)$ a cylindrical structure on $(A,B)$.
@@
    The Drinfeld element is a boundary gauge transformation for $(\psi,J)=({\operatorname{id}},R_{21}^{-1})$.
@@
    The inverse of the balance is a boundary gauge transformation for $(\psi,J)=(\operatorname{Ad}(D),R_{21}^{-1})$.
@@ 3.4. Dual K-matrix
We now describe the second example of a boudary transfer matrix map, which will be the most relevant for applications in quantum affine algebras (see Section 6). It relies on the gauge transformation described in Section 2.4.
@@
Let $A$ be a balanced Hopf algebra, $B\subseteq A$ a right coideal subalgebra, and $(\psi,J,K)$ a cylindrical structure on $(A,B)$. Then, $\widetilde{K}\coloneqq D\cdot K^{-1}$ is a boundary gauge transformation.
@@ 3.5. Generalized dual K-matrices
Immediately we can generalize Proposition 3.4.1 by using a different cylindrical structure extending the same twist pair on $A$, resulting in a larger family of boundary transfer matrices taking values in the given right coideal subalgebra $B$.
@@
Let $A$ be a balanced Hopf algebra, $B,B^{\prime}\subseteq A$ right coideal subalgebras, $(\psi,J,K)$ a cylindrical structure on $(A,B)$ and $(\psi,J,K^{\prime})$ a cylindrical structure on $(A,B^{\prime})$ (with the same twist pair). Then $\widetilde{K}\coloneqq D\cdot(K^{\prime})^{-1}$ is a boundary gauge transformation for $(\psi,J)$.
@@
    As before, the resulting boundary transfer matrix map $\mathscr{T}^{(V)}=\mathop{\sf Tr}_{2}(\overline{\mathbb{K}}_{\bullet,V})$ is invariant under gauge transformations acting simultaneously on both $(\psi,J,K)$ and $(\psi,J,K^{\prime})$.
@@
    Both examples described in Proposition 3.3.3 are special cases of the construction above. Recall that $({\operatorname{id}},R_{21}^{-1},b)$ is a cylindrical structure on $(A,A)$, therefore so is $(\operatorname{Ad}(D),R_{21}^{-1},Db)$. Therefore, if $({\operatorname{id}},R_{21}^{-1},K)$ is a cylindrical structure on $B$, we obtain $\widetilde{K}=Db^{-1}=u$ as a boundary gauge transformation by choosing $K^{\prime}=b$. Similarly, if $({\operatorname{id}},R_{21}^{-1},K)$ is a cylindrical structure on $B$, we obtain $\widetilde{K}=Db^{-1}D^{-1}=b^{-1}$ as a boundary gauge transformation by choosing $K^{\prime}=Db$.
@@
    Let $\widetilde{K}$ be a boundary gauge transformation as described above. We can prove that the image of the corresponding boundary transfer matrix map $\mathscr{T}\colon[{\operatorname{Mod}}_{\sf fd}(A)]\to B$ is contained in $B\cap S(B^{\prime})$. The proof relies essentially on the same arguments used in [Skl88, Prop. 3].
    When $B^{\prime}=B$, as in Section 3.4, this condition may be very restrictive. For instance, in the case of quantum groups of finite type, it implies that $\mathscr{T}$ is rather trivial, see Section 5.8. It is therefore necessary to consider boundary transfer matrix maps with $B^{\prime}\neq B$ in order to get nontrivial elements. For instance, the boundary transfer matrix map associated to the Drinfeld element recovers the morphism constructed by Kolb in [Kol20], which surjects over the center of a quantum symmetric pair subalgebra (see Remark 1).
    In contrast, this trivialization does not occur in the case of quantum affine algebras, as we observe in Section 6.8. $\triangledown$
@@ 3.6. Completions
$\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ is not a bialgebra in general, but it is equipped instead with a cosimplicial structure, see e.g., [Dav07, ATL19]. In particular, the coproduct map $\Delta:A\to A\otimes A$ lifts to a morphism $\Delta:\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})\to\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}^{2})$ given by $\Delta(f)_{V,W}=f_{V\otimes W}$ for all $V,W\in{\mathcal{C}}$. More generally, let $B$ be a right coideal subalgebra of $A$, ${\mathcal{D}}\subseteq{\operatorname{Mod}}(B)$ a module category over ${\mathcal{C}}$, and ${\mathsf{f}}_{{\mathcal{D}}\boxtimes{\mathcal{C}}}:{\mathcal{D}}\boxtimes{\mathcal{C}}\to\operatorname{Vect}$ the forgetful functor given by $(M,V)\mapsto{\mathsf{f}}_{{\mathcal{D}}}(M\otimes V)$. Then, the coproduct $\Delta:B\to B\otimes A$ lifts to a morphism $\Delta:\operatorname{End}({\mathsf{f}}_{{\mathcal{D}}})\to\operatorname{End}({\mathsf{f}}_{{\mathcal{D}}\boxtimes{\mathcal{C}}})$.
@@ 4.1. Kac-Moody algebras [Kac90]
[…]
@@
[…]
@@ 4.2. Drinfeld-Jimbo quantum groups [Jim85, Dri86, Lus94]
where $q_{i}\coloneqq q^{\epsilon_{i}}$, $k_{i}^{\pm 1}\coloneqq k_{\pm\epsilon_{i}h_{i}}$, together with the quantum Serre relations (see e.g., [Lus94, 3.1.1 (e), Cor. 33.1.5]). We consider the Hopf algebra structure on $U_{q}({\mathfrak{g}})$ uniquely determined by the following coproduct formulae
@@ 4.3. Modules over $U_{q}({\mathfrak{g}})$
  • integrable if all $E_{i},F_{i}$ act locally nilpotently; [^7] This implies that $V$ is completely reducible as a (possibly infinite) direct sum of simple finite–dimensional modules over each $U_{q}(\mathfrak{sl}_{2})$-subalgebra generated by $E_{i},k_{i}^{\pm 1},F_{i}$, for each $i\in I$.
  • in category ${\mathcal{O}}$ if the action of $U_{q}{({{\mathfrak{n}}}^{+})}{}$ is locally finite.888 This implies in particular that, for any $v\in V$, $U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}v=0$ for all but finitely many $\beta\in{\mathsf{Q}}^{+}$. Therefore, ${\mathcal{O}}$ coincides with the category ${\mathcal{C}}^{\operatorname{hi}}$ used in [Lus94, Sec. 3.4.7].

### v2

@@ 2.5. Boundary gauge transformation and transfer matrices
Let ${\operatorname{Mod}}_{\sf fd}(A)]$ be the Grothendieck group of the category of finite-dimensional $A$-modules, i.e., the abelian group generated by the isomorphism classes $[V]$ with $V\in{\operatorname{Mod}}_{\sf fd}(A)$ with the relations $[U]+[W]=[V]$ for every short exact sequence $0\to U\to V\to W\to 0$ in ${\operatorname{Mod}}_{\sf fd}(A)$. Since $A$ is a bialgebra, ${\operatorname{Mod}}_{\sf fd}(A)]$ is a ring, with multiplication map given by $V]\cdot[W]=[V\otimes W]$. Moreover, it is commutative, since $A$ is quasitriangular and thus $V\otimes W]=[W\otimes V]$.
Given a cylindrical structure $(\psi,J,K)$ on $(A,B)$, we aim to construct a ring homomorphism from ${\operatorname{Mod}}_{\sf fd}(A)]$ to $B$. Our approach relies on a distinguished family of gauge transformations.
@@
$V]\mapsto\mathscr{T}^{(V)}\coloneqq\mathop{\sf Tr}_{2}\big(\overline{\mathbb{K}}_{\bullet,V}\big)$ (2.22)
where $V\in{\operatorname{Mod}}_{\sf fd}(A)$. In particular, $\mathscr{T}^{(V)},\mathscr{T}^{(W)}]=0$ for all $V,W\in{\operatorname{Mod}}_{\sf fd}(A)$.
@@
Remark 2.5.3.
The existence of a boundary gauge transformation is a property of the twist pair $(\psi,J)$, since it corresponds to the latter being gauge-equivalent with any twist pair $(\overline{\psi},\overline{J})$ satisfying (2.21). Whenever $(\psi,J)$ is part of a cylindrical structure, the choice of the boundary gauge transformation yields the boundary transfer matrix $\mathscr{T}$ in terms of the gauge-transformed tensor K-matrix. $\triangledown$
@@
$=\mathop{\sf Tr}_{2,3}\Big(\big((\overline{J}_{V,W}^{-1})_{23}\cdot(\overline{\mathbb{K}}_{\bullet,W})_{13}\big)^{\mathsf{t}_{3}}\cdot\big(\big(R_{V^{\overline{\psi}},W}\big)_{23}\cdot(\overline{\mathbb{K}}_{\bullet,V})_{12}\big)^{\mathsf{t}_{3}}\Big)\hskip-5.0pt$ (2.25)
@@ 3.1. Balanced and sovereign Hopf algebras
Let $A$ be a quasitriangular Hopf algebra with invertible antipode $S\colon A\to A$. Then, one has
@@
The invertible element $u\coloneqq m((S^{-1}\otimes{\operatorname{id}})(R))$, where $m\colon A\otimes A\to A$ denotes the multiplication map, is called the Drinfeld element. It satisfies the identities [^6] The element $u$ is $u_{2}^{-1}$ in [Dri90].
@@
  • A balance is also a K-matrix. Namely, if $b$ is a balance, then $(\psi,J,K)=({\operatorname{id}},R_{21}^{-1},b)$ is a cylindrical structure on $(A,A)$. By taking $g=b^{-1}$ we see that it is gauge-equivalent to $({\operatorname{id}},R,1)$.
  • A half-balance is an element $h\in A^{\times}$ such that $\Delta(h)=(h\otimes h)\cdot R$ and $h^{2}$ is central (and hence a balance) [KT09, ST09]. Then a cylindrical structure on $(A,A)$ is given by $(\psi,J,K)=(\operatorname{Ad}(h),1\otimes 1,h)$. By taking $g=h^{-1}$, we see that it is gauge-equivalent to $({\operatorname{id}},R,1)$. $\triangledown$
@@ 3.3. Boundary transfer matrix for balanced Hopf algebras
In the case of balanced Hopf algebras, the defining property of a boundary gauge transformation simplifies and becomes algebraic.
@@
Let $g\in A^{\times}$, $(\psi,J)$ a twist pair, and $(\overline{\psi},\overline{J})$ the gauge-transformed twist pair defined in (2.15). Then $g$ is a boundary gauge transformation if
@@
Let $A$ be a balanced Hopf algebra, $B\subseteq A$ a right coideal subalgebra, and $(\psi,J,K)$ a cylindrical structure on $(A,B)$ with tensor K-matrix $\mathbb{K}$}}.
@@
    The Drinfeld element $u$ is a boundary gauge transformation for $(\psi,J)=({\operatorname{id}},R_{21}^{-1})$. Then, $\overline{\mathbb{K}}=(1\otimes u)\cdot{{\mathbb{K}}}$ yields a boundary transfer matrix.
@@
    The inverse of the balance $b^{-1}$ is a boundary gauge transformation for $(\psi,J)=(\operatorname{Ad}(D),R_{21}^{-1})$. Then, $\overline{\mathbb{K}}=(1\otimes b^{-1})\cdot{{\mathbb{K}}}$ yields a boundary transfer matrix.
@@ 3.4. Dual K-matrix
We now describe the second example of a boudary transfer matrix map, which will be the most relevant for applications in quantum affine algebras (see Section 6).
@@
Let $A$ be a balanced Hopf algebra, $B\subseteq A$ a right coideal subalgebra, and $(\psi,J,K)$ a cylindrical structure on $(A,B)$ with tensor K-matrix $\mathbb{K}$}}. Then, $\widetilde{K}\coloneqq D\cdot K^{-1}$ is a boundary gauge transformation and $\overline{\mathbb{K}}=(1\otimes\widetilde{K})\cdot{{\mathbb{K}}}$ yields a boundary transfer matrix.
@@ 3.5. Generalized dual K-matrices
Proposition 3.4.1 admits an immediate generalization.
@@
Let $A$ be a balanced Hopf algebra, $B\subseteq A$ right coideal subalgebras, $(\psi,J,K)$ a cylindrical structure on $(A,B)$ with tensor K-matrix $\mathbb{K}$}}, and $K^{\prime}\in A^{\times}$ such that
$\Delta(K^{\prime})=J^{-1}\cdot K^{\prime}_{2}\cdot R^{\psi}\cdot K^{\prime}_{1}\,.$ (3.12)
Then, $\widetilde{K}\coloneqq D\cdot(K^{\prime})^{-1}$ is a boundary gauge transformation for $(\psi,J)$ and yields a new boundary transfer matrix $\mathscr{T}^{(V)}=\mathop{\sf Tr}_{2}((1\otimes\widetilde{K})\cdot{{\mathbb{K}}})$.
Proof.
The proof is identical to that of Proposition 3.4.1. ∎
@@
    Both examples described in Proposition 3.3.3 are special cases of the construction above. Recall that $({\operatorname{id}},R_{21}^{-1},b)$ is a cylindrical structure on $(A,A)$ and hence so is $(\operatorname{Ad}(D),R_{21}^{-1},Db)$. Therefore, if $({\operatorname{id}},R_{21}^{-1},K)$ is a cylindrical structure on $B$, we obtain $\widetilde{K}=Db^{-1}=u$ or $\widetilde{K}=Db^{-1}D^{-1}=b^{-1}$ as a boundary gauge transformation by choosing $K^{\prime}=b$ or $K^{\prime}=Db$, respectively.
@@
    Whenever $K^{\prime}$ is part of a cylindrical structure of the form $(\psi,J,K^{\prime})$ for some coideal subalgebra $B^{\prime}\subseteq A$, the construction above is reminiscent of Sklyanin’s two-boundary transfer matrix [Skl88], see Section 6.8. As before, the resulting boundary transfer matrix map $\mathscr{T}^{(V)}=\mathop{\sf Tr}_{2}(\overline{\mathbb{K}}_{\bullet,V})$ is invariant under gauge transformations acting simultaneously on both $(\psi,J,K)$ and $(\psi,J,K^{\prime})$.
    Moreover, the image of the corresponding boundary transfer matrix map $\mathscr{T}\colon[{\operatorname{Mod}}_{\sf fd}(A)]\to B$ is contained in $B\cap S(B^{\prime})$. The proof relies essentially on the same arguments used in [Skl88, Prop. 3]. When $B^{\prime}=B$, as in Section 3.4, this condition may be very restrictive. For instance, in the case of quantum groups of finite type, it implies that $\mathscr{T}$ is rather trivial, see Section 5.8. It is therefore necessary to consider boundary transfer matrix maps with $B^{\prime}\neq B$ in order to get nontrivial elements. For instance, the boundary transfer matrix map associated to the Drinfeld element recovers the morphism constructed by Kolb in [Kol20], which surjects over the center of a quantum symmetric pair subalgebra, see Remark  5.9.2 (1). In contrast, this trivialization does not occur in the case of transfer matrices depending on a spectral parameter for quantum affine algebras, as we observe in Section 6.8.
@@
    Although we had to carefully select the cylindrical structure, Proposition 3.4.1 can be seen as the boundary counterpart to the following statement about universal transfer matrices defined in terms of a quasitriangular bialgebra $A$ universal R-matrix $R$. Namely, a ring homomorphism ${\operatorname{Mod}}_{\sf fd}(A)]\to A$ is defined by the assignment
    $V]\mapsto\mathscr{T}^{(V)}_{\sf per}=\mathop{\sf Tr}_{V}R_{\bullet,V}\in A$
    with $R_{\bullet,V}$ also known as (universal) L-operator. This construction finds its origin in the transfer matrices for vertex models and quantum spin chains with periodic boundary conditions, see, e.g., [Ba82], and plays a key role in the definition of q-characters of quantum affine algebras, see [FR99, FM01]. It permits the following generalization:
    $V]\mapsto\mathscr{T}^{(V)}_{{\sf per},a}=\mathop{\sf Tr}_{V}((1\otimes a_{V})\cdot R_{\bullet,V})\in A$
    where $a\in A^{\times}$ is any group-like element. Proposition 3.5.1 can be viewed as the boundary counterpart of this generalization. $\triangledown$
@@ 3.6. Completions
$\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$ is not a bialgebra in general, but it is equipped instead with a cosimplicial structure, see, e.g., [Dav10, ATL19]. In particular, the coproduct map $\Delta:A\to A\otimes A$ lifts to a morphism $\Delta:\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})\to\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}}^{2})$ given by $\Delta(f)_{V,W}=f_{V\otimes W}$ for all $V,W\in{\mathcal{C}}$. More generally, let $B$ be a right coideal subalgebra of $A$, ${\mathcal{D}}\subseteq{\operatorname{Mod}}(B)$ a module category over $\mathcal{C}$}, and ${\mathsf{f}}_{{\mathcal{D}}\boxtimes{\mathcal{C}}}:{\mathcal{D}}\boxtimes{\mathcal{C}}\to\operatorname{Vect}$ the forgetful functor given by $(M,V)\mapsto{\mathsf{f}}_{{\mathcal{D}}}(M\otimes V)$. Then, the coproduct $\Delta:B\to B\otimes A$ lifts to a morphism $\Delta:\operatorname{End}({\mathsf{f}}_{{\mathcal{D}}})\to\operatorname{End}({\mathsf{f}}_{{\mathcal{D}}\boxtimes{\mathcal{C}}})$.
Remark 3.6.2.
In this setting, the gauge transformation induced by a K-matrix yields a map $\overline{\psi}=\operatorname{Ad}(K)\circ\psi$, which is an automorphism of $\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$, but it does not generally preserve $A\subseteq\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$. $\triangledown$
@@ 4.1. Kac-Moody algebras [Kac90]
Let $\left(\cdot,\cdot\right)$ be the nondegenerate invariant bilinear form on $\mathfrak{g}$} depending on a chosen linear complement $\mathfrak{h}$$^{\prime\prime}$ of $\mathfrak{h}$$^{\prime}$ in $\mathfrak{h}$}. The form $\left(\cdot,\cdot\right)$ induces a nondegenerate pairing between the opposite subalgebras in the triangular decomposition ${\mathfrak{g}}\cong{\mathfrak{n}}^{-}\oplus{\mathfrak{h}}\oplus{\mathfrak{n}}^{+}$, where ${\mathfrak{n}}^{+}\coloneqq{\mathbb{C}}\langle\{e_{i}\}_{i\in I}\rangle$, ${\mathfrak{n}}^{-}\coloneqq{\mathbb{C}}\langle\{f_{i}\}_{i\in I}\rangle$, and we obtain a nondegenerate pairing between the subalgebras ${\mathfrak{b}}^{\pm}\coloneqq\langle{\mathfrak{n}}^{\pm},{\mathfrak{h}}\rangle$.
@@
The bilinear form on $\mathfrak{h}$$^{*}$ dual to $\left(\cdot,\cdot\right)$ will be denoted by the same symbol. Note that its restriction to $\mathsf{P}\times\mathsf{P}$ takes values in $\frac{1}{m}{\mathbb{Z}}$ for some $m\in{\mathbb{Z}}_{>0}$.
@@ 4.2. Drinfeld-Jimbo quantum groups [Jim85, Dri86, Lus94]
where $q_{i}\coloneqq q^{\epsilon_{i}}$, $k_{i}^{\pm 1}\coloneqq k_{\pm\epsilon_{i}h_{i}}$, together with the quantum Serre relations (see, e.g., [Lus94, 3.1.1 (e), Cor. 33.1.5]). We consider the Hopf algebra structure on $U_{q}({\mathfrak{g}})$ uniquely determined by the following coproduct formulae
@@ 4.3. Modules over $U_{q}({\mathfrak{g}})$
  • integrable if all $E_{i},F_{i}$ act locally nilpotently; [^7] This implies that $V$ is completely reducible as a (possibly infinite) direct sum of simple finite–dimensional modules over the $U_{q}(\mathfrak{sl}_{2})$-subalgebra generated by $E_{i},k_{i}^{\pm 1},F_{i}$, for each $i\in I$.
  • in category $\mathcal{O}$} if the action of $U_{q}{({{\mathfrak{n}}}^{+})}{}$ is locally finite. [^8] This implies in particular that, for any $v\in V$, $U_{q}{({{\mathfrak{n}}}^{+})}_{\beta}v=0$ for all but finitely many $\beta\in{\mathsf{Q}}^{+}$. Therefore, $\mathcal{O}$} coincides with the category ${\mathcal{C}}^{\operatorname{hi}}$ used in [Lus94, Sec. 3.4.7].